LUCIFER Training & Preparations for Partner Science:
UT 2010 July 01

Observer: DThompson, KRueff, PHibon
Telescope Operator: SAllanson
Instrument Support: TShih
Telescope Support: BCurtis
SW support: JKraus (on call)

Summary

Details

02:59 Afternoon rain and still unstable, delaying opening. Tout=11.6 DewPt=6.4 RH=70.4% Wind ~5m/s Tmirror1=11.4

Closed dome ACQUIRE mode presets from LUCIFER and IRTC pretending to be LUCIFER worked.

Checked LUCIFER UT time and it agrees with TCS computers

04:36 Opening. RH76%, mostly clear.

Need to power on the calibration unit if it is to be used...

Correct initial collimation/pointing:
  • 04:43 Preset to WT10_247 Parallactic/Track
  • 04:50 Preset to BS9164 to collimate, guider showing ~1 arcsec FWHM
  • Initial collimation good, only needed a little focus (-0.3mm)

INAF_Grazian MOS mask if possible LST=16:11 once done with collimating, too late to take a full set.

Test GCS 2x2 binning on guider camera
  • Edit /lbt/tcs/current/etc/GCS/AIP_L.cfg, change guidecam_binning to 2
  • 04:58 Restart GCSL
  • 04:59 Send ACTIVE Preset to BS9164 (same field we were on)

It seems to be working fine! The only visible problem is that the blue circle is still drawn with a 10 pixel radius should be 5 pixels in the 2x2 binning mode. Will leave it in this binned mode to collect some guiding stats. Seeing is fluctuating 0.8-1.4 arcsec. Use these data to check on guiding rms and the earlier set on this star for 1x1 binned data.

05:14 Preset to BS9162 with off-axis guide star. Worked fine, on acquisition star near edge. More guide stats in 2x2 binning.

05:21 Preset to WT10_356 to correct pointing before spectroscopic acquisition. luci#0005

Thompson: HR10 1645+46
  • Acq at 05:26, Sky #0006, Src#0007 change PA to 286.29 in acq script to put brighter source on slit
  • Re-acq at PA=286.29 Sky#8, src#9 DETXY offsets are totally wrong! 32.87,-49.65 should be ~-46,+1 !!
  • Re-acq, ref source at 1388.4, 1328.9, move source to 998, 1338 (offset -46.85, +1.20 rel DETXY) ref source in slit
  • 06:12 HR10_Hspec (10x300s MER), start at EL~73, FWHM on guider ~1.5 arcsec and fluctuating Files 0017-0026
  • Check image of slit in #0016, it appears tilted on detector
  • FWHM about 2 arcsec by end of script
  • Return to ROTANGLE=162.0 EL=70.6
  • specflatH off=0027, on=0028; specarcH Xe off=0029, on=0030, Ar off=0031, on=0032

07:08 humidity 90% and rising, low clouds overhead so we are closing.

09:33 Re-opening. Humidity still high and fluctuating, but trending down. Problem with right shutter door...pins retracted but it is not opening.

Kuhn: HS1946+7658
  • 09:42 Preset to ACT0403 to check pointing
  • 09:45 HS1946HK_acq, files: 0034 sky, 0035 src, 0036 slit, 0037 though slit. Offset 1.56,1.72
  • 09:55 HS1946HK_sci, files: 0038-0041, guider showing ~1 arcsec or a bit below on average
  • 10:18 specflatHK, files: off=0042-0044 on=0045-0047
  • 10:20 specarcHK, files: Xe off=0048 on=0049 Ar off=0050, on=0051

M57 BrGamma 19h +33
  • 10:25 correct pointing nearby 0052
  • 10:33 BrG_18sci_9sky.txt (stop after 18 images collected). Seeing ~0.75 Files: 0053-55
  • 10:46 Switch to GS2 at PA=180 (probe further off field of view) Files 0056 (acq) 0057-0074 (hit Y limit on M1)

Pedani: 21h 0d
  • 11:35 correct pointing nearby 0075
  • 11:37 B3.2_2231+0111 Ks 0076- Collimation poor (twilight) and stopped wfs a couple images into sequence. Stopped script at 5min.

Morning Darks Start with file 0081

Tomorrow:

Co-pointing Model observations (28 points)
  • Change to IRTC authorized on right size (telescope and LBTtools)
  • Correct pointing and initial collimation
  • Start pointing log
  • Preset with IRTC, GUIDE mode gs=0 at PA=0
  • Take images with IRTC and LUCIFER to record source, timestamp

-- DavidThompson - 01 Jul 2010
Topic revision: r6 - 02 Jul 2010, DavidThompson
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback